OpenStudy (anonymous):

Find two consecutive odd numbers such that the sum of the first number and one-third of the second number is equal to fifty.

OpenStudy (radar):

let 2x + 1 be the first odd number let 2x +1 +2 equal next odd number (2x + 3) As given by the problem: 2x + 1 + (2x + 3)/3=50 2x + 1 +(2x)/3 +1 = 50 Multiplying thru by 3 to clear fraction 6x + 3 +2x + 3 = 150 8x = 150 - 6 8x = 144 x= 18 2x + 1 =37 First odd number 2x + 3 = 39 Nex odd number 37 +39/3 = 37 + 13 = 50 confirmed